OpenStudy (anonymous):

x^2+4x-1=0 Solve the equation by taking square roots.

OpenStudy (anonymous):

Use the quadratic formula

OpenStudy (anonymous):

\[ \left\{\left\{x= -2-\sqrt{5}\right\},\left\{x= \sqrt{5}-2\right\}\right\} \]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Everytime i try it doesn't work

OpenStudy (anonymous):

nevermind! i made a slight error! silly me!

OpenStudy (anonymous):

i got 0.24, -4.24

OpenStudy (anonymous):

\[ x^2+4x-1=0\\ a=1\\ b=4\\ c=-1\\ \Delta =b^2-4ac= 16+4=20\\ \sqrt \Delta= \sqrt{20}= 2 \sqrt 5\\ x= \frac{ -b \pm \sqrt \Delta}{2 a}=\frac{-4\pm 2\sqrt{5}}{2}=-2\pm \sqrt 5 \]

OpenStudy (anonymous):

Numerically the roots are \[ {{x = -4.23607}, {x = 0.236068}} \] but you were asked to find the roots using square roots

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thanks